# What is the nature of Little Red Dots and what is not, MIRI SMILES edition

### v1

@@ 4.2 The prospector-SF code
For prospector-SF (Leja et al. 2017, Johnson et al. 2021), we adopt the following assumptions: we use the MIST stellar evolutionary tracks and isochrones (Choi et al., 2016), a Chabrier (2003) IMF, a range in stellar metallicity between -1.0 and 0.19, and gas phase metallicity between -2.0 and 0.5. The ionization parameter for the nebular emission ranges between -4 and -1. The nebular line and continuum emission is generated using the CLOUDY (Ferland et al., 1998). We base the attenuation law on the dust model that combines: (1) the Charlot & Fall (2000) two-component, birth-cloud vs. diffuse dust screens, approach; (2) the Kriek & Conroy (2013a) method that parametrizes the diffuse component as a combination of a Calzetti attenuation plus a Lorentzian Drude to model the strength of the UV bump. Both components of the diffuse dust are then modulated by a power-law factor that varies the slope of the attenuation between -1 to 0.4 relative to Calzetti (). The parameters being fit in this case are the ratio of the nebular to diffuse attenuation, which ranges between 0 and 2, but follows a clipped normal prior centered in 1, and the dust index . For the SFH, we use a non-parametric model, following the flexible SFH prescription (Leja et al., 2019a) with 6 time bins and the bursty-continuity prior (Tacchella et al., 2022). This model includes 5 free parameters that control the ratio of SFR in six adjacent time bins; the first two bins are spaced at Myr and Myr of lookback time, and the remaining four bins are log-spaced to a maximum age of 100 Myr.
@@ 4.3 The prospector-AGN code
A third type of fit, prospector-AGN hereafter, makes use of a hybrid galaxy+AGN model similar to the one described in Barro et al. (2023). Briefly, we use a combination of a stellar emission that dominates in the rest-frame UV and an AGN that dominates in the optical to IR wavelengths. By construction, this is the only model where the AGN makes up the bulk of the luminosity and, as such, it serves as a feasibility test for AGN-dominated SEDs (forced to be dominant, not free as in the case of the prospector-AGN code presented in the next section). The stellar emission is modeled with Prospector using a parametric delayed- SFH with no attenuation. The AGN has two distinct components that account for the emission of the accretion disk and the torus, respectively. The accretion disk is modeled after the empirical QSO templates mentioned above from 0.1 m up to 0.6 m rest-frame, followed by a power-law (f, see for example Hernán-Caballero et al. 2016 or more recently in Bosman et al. 2023 and Killi et al. 2023) with variable spectral index ranging between 0 to 0.5, with default value of , and attenuated with a Calzetti et al. (2000) law. The torus emission is modeled using the clumpy torus models from Nenkova et al. (2008) included in prospector (Leja et al., 2018). These have two free parameters: the optical depth, which ranges between and 150, and the total IR luminosity. Since these parameters are independent of each other, the models allow for AGN-dominated SEDs where the emission from the torus dominates at longer wavelengths ( m), as opposed to the typical QSO templates where the torus emission dominates at m or blueward. The prospector-AGN code discussed next also allows for a similar modeling of the torus emission in the NIR by using empirical templates that account for two separate dust components, warm and hot, peaking at 10m and 50m (see e.g. Lyu et al. 2017 for a discussion of warm- and hot-dust deficient SEDs). The main difference between the AGN modeling codes is in the accretion disk component. prospector-AGN, uses an empirical template with a red NIR slope starting at 1m, similar to other empirical QSO templates, while prospector-AGN uses the aforementioned declining power-law with .
@@ 4.4 The prospector-AGN+ code
Last but not least, prospector-AGN is a modified version of the original prospector code with the Flexible Stellar Population Synthesis (FSPS; Conroy et al. 2009; Conroy & Gunn 2010) for the stellar component and a set of semi-empirical SED models for the AGN component and the galaxy dust emission component, which has been optimized for AGN identification and characterization (Lyu et al., 2022, 2023). For this work, we adopted a similar model configuration as in Lyu et al. (2023) for the SMILES+JADES AGN identification: 1) for the stellar component, we assumed a Kroupa (2001) initial mass function and delayed-tau star-formation history, and adopted the Calzetti attenuation curve with a flexible slope as introduced in Kriek & Conroy (2013b). The stellar nebular line and continuum emission are also turned-on, as pre-configured in FSPS (Byler et al., 2017); 2) the AGN component includes both the AGN-powered continuum from the UV to the far-IR and the narrow and broad emission lines from the UV to the NIR derived based on empirical observations. The continuum shape and line strengths of the AGN SED can be modified by a hybrid extinction configuration featuring the SMC-like curve for the commonly seen UV-optical extinction in Type-1 AGNs and an empirical attenuation law for the AGN IR obscuration; 3) For the galaxy dust emission, since the objects of this work have , we adopted the empirical IR SED model of Haro 11, a low-metallicity star-bursting dwarf galaxy that is believed to share typical features of first-generation galaxies in the early Universe (Lyu et al., 2016; De Rossi et al., 2018). Readers are highly recommended to check Lyu et al. (2023) and references therein for more details.
@@ 5.1 Galaxy by galaxy analysis for the Golden Five LRDs
In this section, we present our analysis of the SEDs of the LRDs detected in several MIRI bands. We concentrate our discussion of properties on the 5 galaxies detected in F1280W and redder bands, which we call the Golden Five sample. At the end of the section, we describe the properties for the galaxies detected up to F1000W and we briefly discuss the rest of the sample, whose fits are presented in Appendix C. The main physical properties of all the LRDs are given in Table 2.
5.1.1 Galaxies detected up to F1800W: JADES-57356 and JADES-204851
Figures 5 and 6 show our fits for JADES57356 and JADES204851, the 2 LRDs in our sample detected up to (at least) 18 m.
@@ 5.1.1 Galaxies detected up to F1800W: JADES-57356 and JADES-204851
The SED fit with synthesizer-AGN uses 2 stellar populations (with independent attenuations), one with a young (60 Myr) and mildly unobscured ( mag) starburst, which also contributes to the faint emission lines, and an older (250 Myr) stellar population with much larger reddening, mag, that dominates the stellar mass content (98% of the total). The fact that no emission lines are observed in the optical ([OIII] or H), but the UV spectral range is flat, favors the contribution of a QSO to that UV emission, outshined by the stars in the optical. From a complementary perspective, very young stars would contribute with emission lines in the optical (as in JADES204851, discussed next), and thus are disfavored by this modeling.
We remark that synthesizer-AGN does use a Calzetti law but assumes independent extinctions for the 2 stellar populations in the host, and also for the AGN, which allows a good fit to the data for JADES57356. Prospector-AGN does not require a two-component fit but it allows for hybrid AGN+galaxy models, which can sometimes fit the SED with 2 distinct components, each dominating a different spectral range, but can also fit the whole SED with only one of those components. This is the case for JADES57356 which, as noted above, exhibits a similar best-fit model to the prospector-SF model for the UV-to-NIR region.
While all the models obtain similarly good UV-to-NIR fits, they differ substantially in the MIR emission, where the SED shows a clear flattening beyond 1 m (F770W-F1000W0 mag) followed by an upturn around 2m that steepens quickly towards the redder bands. This SED shape is well reproduced by the stellar-continuum-dominated models that combine a stellar peak around 1.6m with dust emission presenting a prominent 3 m PAH line in the MIRI F2100W band. For synthesizer-AGN, the overall best-fitting code ( values given in the plots), the dust emission model is quite extreme: the best template corresponds to a ULIRG with mag, 350 pc star-forming region size, 60% luminosity ratio of OB stars in hot spots compared to total luminosity, and dust density in hot spots cm. This dust emission model peaks at rest-frame m (see Figure 16 in Appendix B), quite a blue wavelength compared to the more typical m for (U)LIRGs (e.g., Rieke et al., 2009), revealing the important role of warm/hot dust in LRDs, and also implying a relatively low emission at (sub-)mm wavelengths, where the flux of LRDs is faint (Labbé et al., 2023b; Williams et al., 2023a).
JADES204851, whose SED and postage stamps are shown in Figure 6, is a source with spectroscopic redshift discussed previously in Matthee et al. (2023b, GOODS-S-13971 in that paper), where they reported on the presence of a broad 2200 km s H component (S/N5 and with some artifacts in the spectrum –certainly, this is not one of the clearest BLR sources in Matthee et al. 2023b–) arising from an AGN with a M SMBH. This AGN would account for half of the H emission, according to the spectroscopic analysis.
In our synthesizer-AGN fits for JADES204851 (top left panel of Figure 6), the AGN emits around 10% of the total emission at 2 m, also contributing to emission lines such as H, [OIII], and MgII. The MIR emission would also have a significant contribution from star-formation powered dust emission, with a % possible contribution from a QSO-like emission. For this galaxy, the best fitting dust emission model corresponds to one 350 pc ULIRG star-forming region with highly embedded stars ( mag), with a 40% ratio of OB stars in hot spots compared to the total luminosity, and dust density in hot spots cm. The optical-to-NIR is fitted with a combination of two stellar populations. One of them is a very young (1 Myr) unobscured ( mag) starburst, which takes care of the strong emission seen in spectroscopy and detected in the NIRCam imaging and the blue continuum. The other population is slightly older (10 Myr) affected by a large reddening ( mag) and dominating (94%) the total stellar mass ().
5.1.2 Galaxy detected up to F1500W: JADES-219000
Figure 7 discusses the results for JADES219000, the third LRD in our sample detected up to 15 m. This source is presented in Sun et al. (2024, in prep.), with a spectroscopic redshift of . Similarly to JADES57356, the best-fit with synthesizer-AGN indicates a UV-to-NIR SED dominated by stars with a relatively high mass (), and dust emission revealing a very intense radiation field. The dust emission model corresponds to an intense starburst with 90% luminosity arising from OB stars embedded in a mag compact (350 pc), dense ( cm) and clumpy dust cloud. The very hot dust present in this object could also be heated by an AGN, but the typical torus emission used in prospector-AGN fails to fall within the F1800W and F2100W upper limits. This, in part, explains the larger value compared to the other codes, jointly with discrepancies in the FUV which could be interpreted as attenuation law effects.
5.1.3 Galaxies detected up to F1280W: JADES-211388 and JADES-79803
Figure 8 shows the SED fits for JADES211388, a spectroscopically confirmed LRD at . This is one of the highest redshift LRDs and, consequently, the MIRI detections probe shorter rest-frame wavelengths than in previous galaxies.
5.2 Summary and implications for the nature of the NIR emission of the LRDs
@@
Overall, all 4 models fit the characteristic bimodal SED of the Golden Five LRDs relatively well. Qualitatively, the UV-optical spectral range is dominated either by 1) stars: two young populations with very different attenuations in synthesizer-AGN, or a single population with an extremely gray attenuation in prospector-SF and prospector-AGN; or 2) emission from an obscured accretion disk combined with a young stellar population that contributes only in the UV. In the next section, we discuss the implied stellar masses and other stellar properties for the Golden Five and all other LRDs that have very similar UV-to-optical SEDs (see, e.g., Figure 11) probed primarily by the NIRCam bands. However, the Golden Five, and other MIRI-detected subsamples, allow us to probe further into the rest-frame NIR of the LRDs to characterize the origin of the emission in that spectral range.
@@
Previous LRD papers fitting NIRCam-only SEDs with empirical AGN templates (e.g., Kocevski et al. 2023; Barro et al. 2023; Greene et al. 2023) implied that the steep rest-optical slope would continue into the NIR. This trend seemed to be confirmed by the handful of LRDs with MIRI detections up to F770W and F1000W (Akins et al. 2023; Barro et al. 2023). However, as shown in section 3.1 and also Williams et al. (2023a), the MIRI data at longer wavelengths indicate that LRDs have a flattening in the SED between 1 and 2 m. Interestingly, the SEDs of 4 out of the Golden Five LRDs at redshifts , which have direct detections of the NIR continuum at 2 m, appear to show upturns with different slopes at longer wavelengths. This suggests that, while it does not dominate, the amount of dust emission from star formation or an AGN can vary substantially from galaxy to galaxy.
@@
Our 4 codes fit the NIR spectral range relatively well with a combination of the dominant source of UV-optical emission (i.e., stars or an accretion disk) plus a variable contribution from dust emission, either from star-formation (synthesizer-AGN, and prospector-SF) or from an AGN torus (prospector-AGN, and prospector-AGN can do either depending on the source). The left panel of Figure 10 illustrates this variation showing the rest-frame color-color diagram for the LRDs and some templates and models. The 0.4-to-1 m color probes the optical to NIR slope. This color is similar to the F277W-F444W used in the sample selection but it is not affected by emission lines and, thus, is a better proxy for the amount of dust attenuation. The color is also similar to the rest-frame which is a known tracer of large dust attenuation in the UVJ diagram ( mag for very dusty galaxies, e.g., Brammer et al. 2011, Wuyts et al. 2011). The 1-to-3 m color probes the NIR slope which tracks the deviation from a pure stellar continuum due to the contribution of dust and also nebular continuum emission, as shown by the templates in the right panel. The black horizontal line shows the flat 1-to-3 m color of a stellar population with zero contribution from dust emission, which peaks at 1.6 m (solid black line in the right panel). The magenta region indicates the redder colors up to 0.5 mag relative to the stellar continuum due to increasing amounts of nebular continuum (magenta dashed line in the right panel). The 3 grey lines indicate the color tracks with increasing attenuation (A(V)=1 to 4) for the Polletta et al. (2007) QSO1 template (solid), the hot dust deficient (HDD) template of Lyu et al. (2017) used in prospector-AGN+ (dashed-dotted), and the accretion disk model with declining slope () and zero dust emission used in prospector-AGN (dashed).
@@
The green and blue lines in both panels of Figure 10 illustrate how the modeling codes populate the color-color diagram in between the flat (color mag) stellar-only sequence, and the hot-dust dominated sequence of the QSO1 template with increasing contributions from dust emission (i.e., larger IR-luminosities) from star-formation or an AGN, relative to the stellar or accretion disk continua. The color-color tracks ranging from moderate to high infrared luminosities at the top is computed by scaling the dust emission (f/f5% to 60% at 2 m) of JADES57356. As expected, the star-forming dust templates exhibit larger luminosities than the torus at similar 1-to-3 m colors because their SEDs extend to longer wavelengths with a more prominent peak (see also Figure 16). While a single NIR color is not able to capture all the nuances of the different dust emission templates, Figure 16 in the appendix shows that the torus and star-forming dust templates in Siebenmorgen & Krügel (2007) and Nenkova et al. (2008) have similar slopes for the same normalization (i.e., same f/f value). Consequently, the four models modeling codes can, in principle, reproduce the NIR continuum colors of the LRDs. However, to do so with purely AGN-dominated model requires an accretion disk model with declining slope (dashed grey line) and only a small contribution from dust torus emission (thin blue line) to successfully reproduce the bluest 1-to-3 m1.5 colors at the lower limit of the HDD template. Note, that prospector-AGN+ can still reproduce those colors with a hybrid of stellar and AGN emission. The main difference between the star-formation and torus dust-emission models, and the best chance to precisely identify the source of the NIR emission, is the presence of the PAH emission line in the star-formation models at 3 m, which is hinted in JADES57356.
@@
The 1-to-3 m colors of the LRDs (circles for direct detections and triangles for upper limits) are bracketed between the stellar and QSO sequences, but the Golden Five galaxies with direct detections beyond rest-frame 2 m exhibit colors that are at least 1 mag redder than a flat, stellar-only SED. JADES57356 and JADES79803 in particular have redder colors, [1-to-3 m]2 mag, indicative of larger dust emissions and IR luminosities. JADES-204851 and JADES-219000 have bluer colors, [1-to-3 m]1 mag. The latter and JADES-211388 are not well constrained beyond 2 m because of their higher redshifts, but upper/lower triangles show the range in possible colors spanned between the prospector-SF and synthesizer-AGN best-fits, which feature different amounts of dust emission within the upper limits of the redder MIRI bands. The broad range highlights the need for deep, long wavelength MIRI data to constrain precisely the amount and heating nature of dust emission in LRDs.
@@
This issue is exacerbated for the LRDs detected only at shorter wavelengths. Figures 11 and 17 show the stacked SEDs for the LRDs detected up to F1000W and F770W, which reveal that their SEDs are only well constrained up to rest-frame 1.6 m and 1 m, respectively. Consequently, the 1-to-3 m colors of their best-fit models span a much larger range from the synthesizer-AGN, stellar-only fits with [1-to-3 m]0 mag (e.g., top-left panel of Figure 11), to the much redder best-fit models of prospector-AGN+ and prospector-AGN (bottom-left and right), which sometimes fit the MIRI upper limits with pronounced upturns at 2 m. It is worth mentioning the F770W-only sample places more restrictive constraints against very red QSO-like colors than the F1000W-only sample. This is because, as shown in Figure 4, the F1000W-only LRDs have redder F777W-F1000W colors than the upper limits of the F770W-only LRDs and thus lead to a stronger flattening of the SED in the 1-2m range. Consequently, the possible upturn after 2 m is not nearly as red.
@@
In summary, the exact contribution from dust emission to the NIR SEDs of the LRDs is still poorly constrained by the available MIRI data at long wavelengths for a significant fraction of our full LRD sample. Nonetheless, the constraints from the Golden Five sample point to a certain diversity in the dust emission. This emission is lower than prior expectations based on QSO templates but is clearly larger than stellar-only SEDs. Overall, concerning dust, the main conclusion is that most LRDs could harbor some (relatively large) amount of dust emission and the heating source must be intense, although not necessarily a dominant obscured AGN for all sources.

### v2

@@ 4.2 The prospector-SF code
For prospector-SF (Leja et al. 2017, Johnson et al. 2021), we use the MIST stellar evolutionary tracks and isochrones (Choi et al., 2016), a Chabrier (2003) IMF, a range in stellar metallicity between -1.0 and 0.19, and gas phase metallicity between -2.0 and 0.5. For the SFH, we use a non-parametric model, following the flexible SFH prescription (Leja et al., 2019a) with 6 time bins and the bursty-continuity prior (Tacchella et al., 2022). The ionization parameter for the nebular emission ranges from = -4 to -1. The nebular line and continuum emission are generated using CLOUDY (Ferland et al., 1998). We base the attenuation law on a dust model that combines: (1) the Charlot & Fall (2000) two-component approach, birth-cloud vs. diffuse dust screens; and (2) the Kriek & Conroy (2013a) method that parametrizes the diffuse component as a combination of a Calzetti attenuation plus a Lorentzian Drude to model the strength of the UV bump. Both components of the diffuse dust are then modulated by a power-law factor that varies the slope of the attenuation between -1 to 0.4 relative to Calzetti ().
This model includes 5 free parameters that control the ratio of SFR in six adjacent time bins; the first two bins are spaced at Myr and Myr of lookback time, and the remaining four bins are log-spaced to a maximum age of 100 Myr. In addition, it fits (1) for the ratio of the nebular to diffuse attenuation, which ranges between 0 and 2, but follows a clipped normal prior centered on 1, and also (2) for the dust index .
@@ 4.3 The prospector-AGN code
A third type of fit, prospector-AGN hereafter, makes use of a hybrid galaxy+AGN model similar to the one described in Barro et al. (2023). Briefly, we use a combination of a stellar emission component that dominates in the rest-frame UV and an AGN one that dominates in the optical to IR wavelengths. By construction, this is the only model where the AGN makes up the bulk of the luminosity and, as such, it serves as a feasibility test for AGN-dominated SEDs (forced to be dominant, not free as in the case of the prospector-AGN code presented below). The stellar emission is modeled with Prospector using a parametric delayed- SFH with no attenuation. The AGN has two distinct components that account for the emission of the accretion disk and the torus, respectively. The accretion disk is modeled after the empirical QSO templates mentioned above from 0.1 m up to 0.6 m rest-frame, followed by a power-law (f, see for example Hernán-Caballero et al. 2016 or more recently in Bosman et al. 2023 and Killi et al. 2023) with variable spectral index ranging between 0 to 0.5, with default value of , and attenuated with a Calzetti et al. (2000) law. The torus emission is modeled using the clumpy torus models from Nenkova et al. (2008) included in prospector (Leja et al., 2018). These have two free parameters: the optical depth, which ranges between and 150, and the total IR luminosity. Since these parameters are independent of each other, the models allow for AGN-dominated SEDs where the emission from the torus dominates at longer wavelengths ( m), as opposed to the typical QSO templates where the torus emission dominates at m or blueward.
@@ 4.4 The prospector-AGN+ code
Last but not least, prospector-AGN is a modified version of the original prospector code with the Flexible Stellar Population Synthesis (FSPS; Conroy et al. 2009; Conroy & Gunn 2010) for the stellar component. We assumed a Kroupa (2001) initial mass function and delayed-tau star-formation history. The stellar nebular line and continuum emission are also turned-on, as pre-configured in FSPS (Byler et al., 2017). We adopted the Calzetti attenuation curve with a flexible slope as introduced in Kriek & Conroy (2013b). For the galaxy dust emission, since the objects of this work have , we adopted the empirical IR SED model of Haro 11, a low-metallicity star-bursting dwarf galaxy that is believed to share typical features of first-generation galaxies in the early Universe (Lyu et al., 2016; De Rossi et al., 2018).
This code uses a set of semi-empirical AGN SED models, which have been optimized for AGN identification and characterization (Lyu et al., 2022, 2023). For this work, we adopted a similar model configuration as in Lyu et al. (2023) for the SMILES+JADES AGN identification: the AGN component includes both the AGN-powered continuum from the UV to the far-IR and the narrow and broad emission lines from the UV to the NIR derived based on empirical observations. The continuum shape and line strengths of the AGN SED can be modified by a hybrid extinction configuration featuring the SMC-like curve for the commonly seen UV-optical extinction in Type-1 AGNs and an empirical attenuation law for the IR obscuration. Further details can be found in Lyu et al. (2023) and references therein.
In total, this code uses 7 free parameters to describe the stellar component and 2 for the AGN.
4.5 Summary
These four codes let us compare models of the LRDs with a wide variety of input assumptions. This is important to let us deduce which properties are likely to be intrinsic rather than due to modeling issues. synthesizer-AGN, prospector-SF, and prospector-AGN+ use three different prescriptions for the stellar populations, with three varieties for the interstellar extinction law. synthesizer-AGN and prospector-AGN+ treat the stellar-heated dust emission in two different ways. synthesizer-AGN, prospector-AGN, and prospector-AGN+ use three different models for the AGN emission. This latter variety is particularly important given the diversity of possibilities for AGN SEDs and the importance of capturing this variety in modeling LRDs. The first case, synthesizer-AGN, uses an empirical fixed model SED. prospector-AGN fits empirical templates obscured by an attenuation law derived specifically for AGNs (see e.g. Lyu et al. 2017), while prospector-AGN uses a declining power-law with in place of the empirical templates.
@@ 5.1 Galaxy by galaxy analysis for the Golden Five LRDs
In this section, we present our analysis of the SEDs of the LRDs detected in several MIRI bands. We concentrate our discussion of properties on the five galaxies detected in F1280W and redder bands, which we call the Golden Five sample. We compare the global results of the SED fitting provided by each one of our 4 different codes in different subsections, and comparatively discuss how they fit the UV, optical and near-infrared spectral regions. At the end of the section, we describe the properties for the galaxies detected up to F1000W and we briefly discuss the rest of the sample. A more detailed discussion of all the galaxies in our our sample is presented in Appendix C. The main physical properties of all the LRDs are given in Table 2.
5.1.1 Results for the Golden Five galaxies with synthesizer-AGN
@@
With synthesizer-AGN, the rest-frame UV spectral range is fitted with a young stellar population for all the galaxies in the Golden Five sample except for JADES57356, which is fitted with the QSO component. For this galaxy, the fact that no emission lines are observed in the optical ([OIII] or H), but the UV spectral range is flat, favors the contribution of a QSO to the UV emission, outshined by the stars in the optical. Very young stars would contribute strong emission lines in the optical (as in JADES204851), and thus are disfavored by this modeling. The different UV (relatively flat) slopes observed for the Golden Five sample are reproduced with 1 Myr old (mass-weighted age) stellar population and mag dust attenuation, which is also able to reproduce the emission lines observed for four of the Golden Five galaxies.
The rest-frame optical range of the Golden Five sample is dominated by slightly more evolved stars in all cases, typically of 10-100 Myr mass-weighted age. They account for most of the stellar mass (typically, more then 90%) but do not contribute much, in relative terms with respect to the younger stars, to the optical emission lines. This more evolved stellar populations present high attenuations, typically mag.
If we consider the ratio of the attenuation of the far-UV with respect the optical emission arising from all the stellar populations included in the synthesizer-AGN modeling, , we typically find flatter slopes, , than what is expected for the Calzetti et al. (2000) attenuation law (). This effect is, in fact, common to all modeling techniques, they all typically obtain gray attenuation laws.
Finally, the rest-frame near/mid-infrared emission of the Golden Five LRDs is fitted with dust combined with star formation in three Golden Five galaxies, and dust in an AGN torus in the other two. The heating source is indicated by the possible detection of a most probably star-formation related 3 m PAH feature in JADES57356, and by the differences in slopes of the SEDs for the other sources, steeper in the case of AGN-dominated fits (JADES211388 and JADES79803).
The synthesizer-AGN code presents the smallest values for the fits of the SEDs of all the Golden Five galaxies except one.
5.1.2 Results for the Golden Five galaxies with prospector-SF
By construction, prospector-SF only considers stars and dust heated by stars to fit the SEDs. In general, the values obtained by this code and the following one are larger than those obtained with synthesizer-AGN (except for JADES219000) and smaller than prospector-AGN.
The prospector-SF results imply a 30 Myr stellar population (mass-weighted age), with large attenuations mag and a flat attenuation law. These properties translate to larger stellar masses compared to other codes (a 0.5 dex difference with respect to synthesizer-AGN, for example). We remark that the dust emission models are significantly colder than what is obtained with synthesizer-AGN (see Appendix B).
5.1.3 Results for the Golden Five galaxies with prospector-AGN
The main distinct characteristic of the results achieved with prospector-AGN for the Golden Five galaxies is the smaller stellar mass. Given that, by construction, this code only fits the rest-frame UV spectral region with stars, and the optical and infrared with AGN templates (including emission from the accretion disk and the torus), the stellar masses are up to a factor of 100 smaller than those estimated with the other codes. In general, better fits in terms of values are obtained with prospector-AGN compared to prospector-SF, except for JADES57356, the galaxy with possible PAH emission. But better fits are obtained with the mixed models in synthesizer-AGN (favoring stellar emission in the UV and optical, a variety of results in the near-infrared).
5.1.4 Results for the Golden Five galaxies with prospector-AGN+
The prospector-AGN code fits to the SEDs of the Golden Five galaxies are very similar to those outlined for synthesizer-AGN. The rest-frame UV and optical spectral ranges are dominated by stars, except at 0.8-1.0 m for two galaxies (JADES211388 and JADES79803) whose infrared emission was fitted with an AGN torus model. This was also the best solution for synthesizer-AGN. The only significant difference with respect to the results achieved with synthesizer-AGN is found for the infrared emission of JADES219000, whose slope if we take the MIRI upper limits at face value is too steep for a star formation model, so prospector-AGN (which used upper limits as regular points) prefers a torus template (while the synthesizer-AGN star-formation heated dust model lies below the upper limits). The typical mass-weighted ages for the fits with the prospector-AGN code are around 150 Myr, with extinctions around mag with a flatter attenuation law compared to the Calzetti et al. (2000) recipe.
A detailed analysis of the SED fits for each galaxy in the Golden Five sample is presented in Appendix C.
5.2 Implications for the nature of the NIR emission of the LRDs
@@
Overall, all four models fit the characteristic bimodal SED of the Golden Five LRDs relatively well. Qualitatively, the UV-optical spectral range is dominated either by 1) stars: two young populations with very different attenuations in synthesizer-AGN, or a single population with an extremely gray attenuation in prospector-SF and prospector-AGN; or 2) emission from an obscured accretion disk combined with a young stellar population that contributes only in the UV. In the next section, we discuss the implied stellar masses and other stellar properties for the Golden Five and all other LRDs that have very similar UV-to-optical SEDs (see, e.g., Figure 11) probed primarily by the NIRCam bands. However, the Golden Five, and other MIRI-detected subsamples, allow us to probe further into the rest-frame NIR of the LRDs to characterize the origin of the emission in that spectral range.
@@
Our results confirm the flattening of the LRD SEDs in the rest near infrared. Previous LRD papers fitting NIRCam-only SEDs with empirical AGN templates (e.g., Kocevski et al. 2023; Barro et al. 2023; Greene et al. 2023) implied that the steep rest-optical slope would continue into the NIR. This trend seemed to be confirmed by the handful of LRDs with MIRI detections up to F770W and F1000W (Akins et al. 2023; Barro et al. 2023). However, as shown in section 3.1 and also Williams et al. (2023a), the MIRI data at longer wavelengths indicate that LRDs have a flattening in the SED between 1 and 2 m (rest). Interestingly, the SEDs of four out of the Golden Five LRDs at redshifts , which have direct detections of the NIR continuum at 2 m, appear to show upturns with different slopes at longer wavelengths. This suggests that, while it does not dominate, the amount of dust emission from star formation or an AGN can vary substantially from object to object.
@@
Our four codes fit the NIR spectral range with a combination of the dominant source of UV-optical emission (i.e., stars or an accretion disk) plus a variable contribution from dust emission, either from star-formation (synthesizer-AGN, and prospector-SF) or from an AGN torus (prospector-AGN, and prospector-AGN depending on the source). The left panel of Figure 10 illustrates this variation showing the rest-frame color-color diagram for the LRDs and some templates and models. The 0.4-to-1 m color probes the optical to NIR slope. This color is similar to the F277W-F444W used in the sample selection but it is not affected by emission lines and, thus, is a better proxy for the amount of dust attenuation. The color is also similar to the rest-frame which is a known tracer of large dust attenuation in the UVJ diagram ( mag for very dusty galaxies, e.g., Brammer et al. 2011, Wuyts et al. 2011).
The behavior can be discussed in terms of the color-color behavior in Figure 10. The black horizontal line in the figure shows the flat 1-to-3 m color of a stellar population with zero contribution from dust emission, which peaks at 1.6 m. The magenta region indicates the redder colors up to 0.5 mag relative to the stellar continuum due to increasing amounts of nebular continuum (magenta dashed line in the left panel). The 3 grey lines indicate the color tracks with increasing attenuation (A(V)=1 to 4 mag) for the Polletta et al. (2007) QSO1 template (solid), the hot dust deficient (HDD) template of Lyu et al. (2017) used in prospector-AGN+ (dashed-dotted), and the accretion disk model with declining slope () and zero dust emission used in prospector-AGN. The right panel of Figure 10 illustrates some of same trends as they affect the SED templates.
@@
Figure 10 illustrates how the modeling codes populate the color-color diagram between the flat (color mag) stellar-only sequence, and the hot-dust dominated sequence of the QSO1 template with increasing contributions from dust emission (i.e., larger IR-luminosities) from star-formation or an AGN, relative to the stellar or accretion disk continua. The color-color tracks ranging from moderate to high infrared luminosities (indicated in log for the star-forming case and log for AGNs) are computed by scaling the dust emission (fdust/ftotal5% to 60% at 2 m) of JADES57356. As expected, the star-forming dust templates exhibit larger luminosities than the torus at similar 1-to-3 m colors because their SEDs extend to longer wavelengths with a more prominent peak (see also Figure 16 in the appendix). While a single NIR color is not able to capture all the nuances of the different dust emission templates, Figure 16 shows that the torus and star-forming dust templates in Siebenmorgen & Krügel (2007) and Nenkova et al. (2008) have similar slopes for the same normalization (i.e., same fdust/ftotal value). Consequently, the four modeling codes can all, in principle, reproduce the NIR continuum of the LRDs.
@@
The Golden Five galaxies with direct detections beyond rest-frame 2 m exhibit colors that are at least 1 mag redder than a flat, stellar-only SED. In particular, two of them (JADES57356 and JADES79803) have very red colors, [1-to-3 m] 2 mag, indicative of large dust emissions and IR luminosities. JADES-219000 and JADES-211388 are not well constrained beyond 2 m because of their higher redshifts, but upper/lower triangles show the range in possible colors spanned between the prospector-SF and synthesizer-AGN best-fits, which feature different amounts of dust emission within the upper limits of the redder MIRI bands. These limits fall within the overall behavior of all the sources in the figure. The broad range overall in Figure 10 highlights the need for deep, long wavelength MIRI data to constrain precisely the amount and heating nature of dust emission in LRDs.
However, we can already see that fitting the colors with a purely AGN-dominated model requires an accretion disk model with declining slope (dashed grey line) and only a small contribution from dust torus emission (thin blue line) to successfully reproduce the bluest 1-to-3 m1.5 colors at the lower limit of the HDD template. This would differ significantly from lower redshift AGN, which tend to have strong emission from their circumnuclear tori. This issue is mitigated by prospector-AGN+, which can reproduce those colors with a hybrid of AGN and stellar emission.
@@
We now discuss the LRDs detected only at shorter wavelengths. Figures 11 and 17 show the stacked SEDs for the LRDs detected up to F1000W and F770W, which reveal that their SEDs are only well constrained up to rest-frame 1.6 m and 1 m, respectively. Consequently, the 1-to-3 m colors of their best-fit models span a much larger range from the synthesizer-AGN, stellar-only fits with [1-to-3 m] 0 mag (e.g., top-left panel of Figure 11), to the much redder best-fit models of prospector-AGN+ and prospector-AGN (bottom-left and right), which sometimes fit the MIRI upper limits with pronounced upturns at 2 m. It is worth mentioning the F770W-only sample places more restrictive constraints against very red QSO-like colors than the F1000W-only sample. This is because, as shown in Figure 4, the F1000W-only LRDs have redder F777W-F1000W colors than the upper limits of the F770W-only LRDs and thus lead to a stronger flattening of the SED in the 1-2m range. Consequently, the possible upturn after 2 m is not nearly as red.
@@
In summary, the exact contribution from dust emission to the NIR SEDs of the LRDs is still poorly constrained by the available MIRI data at long wavelengths. Nonetheless, the constraints point to a certain diversity in the dust emission. This emission is clearly larger than stellar-only SEDs but in many cases lower than prior expectations based on QSO templates. Overall, the main conclusion is that most LRDs could harbor some (relatively large) amount of dust emission and the heating source must be intense, although not necessarily a dominant obscured AGN.